26. A high school band teacher has a record of each student's absence. The results, in days, are summarized below.

Days

Frequency

0

9

1

11

2

5

3

4

4

2

5

1

6

1

7

1

      The relative frequency of the students with at most 3 absences equals:

         [A]     0.265          [B]     0.853          [C]     0.735          [D]    0.882         
Answer #1

Solution :

Given that

Total = 9 + 11 + 5 + 4 + 2 + 1 + 1 + 1 = 34

The relative frequency of the students with at most 3 absences equals

= P(X 3)

= (9 + 11 + 5 + 4 ) / 34

= 29 / 34

=0.853